determine whether the following equations have solutions that are two different rational solutions:two different irrational solutions:exactly one rational solution or two different imaginary solutions.
3 + 8z^2= -7z

Answer by stanbon(75887)

determine whether the following equations have solutions that are two different rational solutions:two different irrational solutions:exactly one rational solution or two different imaginary solutions.
3 + 8z^2= -7z
---
Rearrange:
8z^2+7z+3 = 0
a=8 ; b=7 ; c=3
-------------------
Discriminant = b^2-4ac = 49 -4*8*3 < 0
---
Since the discriminant is < 0
the quadrated has two different imaginary solutions.
-----
Cheers,
Stan H.